# Stability of ranks under field extensions

### v3

@@ Abstract.
1. Introduction
Given a tensor $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we denote by $\operatorname{PR}_{\mathbb{K}}(T)$ (resp. $\operatorname{SR}_{\mathbb{K}}(T)$, $\operatorname{AR}_{\mathbb{K}}(T)$, $\operatorname{GR}_{\mathbb{K}}(T)$) the partition rank [^1] It is also called the Schmidt rank in number theory [37]. [36] (resp. slice rank [43], analytic rank [21], geometric rank [28]) of $T$ (cf.  (3)–(6)). These ranks were first introduced to study various problems in additive combinatorics [22, 23, 21, 43, 13, 16, 36], but very soon they also received great attention from other fields such as algebraic geometry [27, 18], coding theory [26, 4] and algebraic complexity theory [5, 28]. Arguably, the most fundamental problems concerning these ranks are their relations to each other and their stability under field extensions. The stability of a function $f$ of tensors under field extensions means the existence of two constants $c\coloneqq c(d,\mathbb{K})>0$ and $C\coloneqq C(d,\mathbb{K})>0$ such that for any $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and any (finite) algebraic extension $\mathbb{F}/\mathbb{K}$, we have
@@ 1. Introduction
where $T^{\mathbb{F}}$ is the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ determined by $T$. Here and in the rest of the paper, we use $C(d,\mathbb{K})$ to indicate the dependence of the constant $C$ on $d$ and $\mathbb{K}$. We remark that in the context of algebraic geometry [24, 12, 31], the stability of $f$ may be rephrased as the invariance of the order of $f$ under base changes.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
We remark that the original formulations of Conjectures 1.11.3 actually require the constant $C$ to be only dependent on $d$. However, for instance, it was proved in [9] that the original version of Conjecture 1.1 is true over finite fields of sufficiently large cardinality depending on the analytic rank. Thus, to completely resolve the original version of Conjecture 1.1, one needs to remove the dependence of the result in [9] on the analytic rank and prove Conjecture 1.1 for fields of small cardinalities. We also want to mention that according to [42, Remark 4], algebraic problems over fields of small cardinalities are usually considered more difficult than those over sufficiently large fields.
@@
According to [1], one can prove that Conjecture 1.3 implies Conjecture 1.1 with an extra assumption. Only very recently, it was shown that this assumption can actually be removed [9]. In Theorem 5.6, we will prove that Conjectures 1.11.3 are all equivalent. In fact, Theorem 5.6 proves that Conjectures 1.11.3 are all equivalent to
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
It is noticeable that Conjecture 1.4 is a weaker version of the direct sum conjecture for partition rank [20], which is an analogue of the direct sum conjecture for cp-rank [41, 25, 29], disproved by [38]. In particular, Conjecture 1.4 is true for slice rank [20]. Thus, Theorem 5.6 trivially implies that Conjectures 1.11.3 are true for $d=3$, which were already known by results in [8, 1, 14].
One of the main ingredients in the proof of Theorem 5.6 is the stability of the analytic rank, proved in Theorem 4.2. In earlier works such as [34, Proposition 8.2] and [9, Proof of Corollary 1], it was proved that for any tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and integers $l\geq 1,d\geq 2$,
$\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)\leq l^{d-1}\operatorname{AR}_{\mathbb{F}_{q}}(T).$
Theorem 4.2 not only improves the coefficient $l^{d-1}$ to a constant depending only on $d$ and $q$, but it also shows that $\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)$ is bounded below by $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, up to a constant independent to $l$. As an interesting application of Theorem 4.2, we prove in Proposition 4.4 that $\operatorname{AR}_{\mathbb{F}_{q}}(T)=\operatorname{GR}_{\mathbb{F}_{q}}(T)$ up to a constant factor. Previously, it was only known that $\operatorname{GR}_{\mathbb{F}_{q}}(T)$ is bounded above by a constant multiple of $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, see [8, Theorem 1] and [1, Proof of Theorem 1.13 (1)]. Moreover, we show that $\lim_{l\to\infty}\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)=\operatorname{GR}_{\mathbb{F}_{q}}(T)$ in Theorem 4.3, which is an analogue of [28, Theorem 8.1].
As a corollary of the proof of Theorem 5.6, we obtain Proposition 5.3 concerning the stability of slice rank. Although a stronger stability result [14] may be obtained by geometric invariant theory, our proof of Proposition 5.3 is much more elementary. The slice rank $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ (cf.  (22)) of a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ was introduced in [1, Definition 1.3] and it was conjectured that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is also stable under field extensions.
@@
[1, Conjecture 1.16] There is a constant $C\coloneqq C(d)>0$ such that for any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq C\operatorname{SR}_{\overline{\mathbb{K}}}(W\otimes\overline{\mathbb{K}}).$
@@
To conclude this section, we summarize the main results of this paper in Figure 1, where $T$ is a tensor and $\mathsf{W}$ is a linear subspace of tensors, $\mathbb{F}$ is a finite extension of $\mathbb{F}_{q}$. The symbol $\approx$” means equal up to a constant factor, “?absent?$\approx$?≈ ?” means $\approx$” is conjectured to hold and all conjectures around $\operatorname{PR}_{\mathbb{F}_{q}}(T)$ are equivalent.
@@
Figure 1. Summary of results
2. Notations
We first recall three basic operations. Given a permutation $\pi\in\mathfrak{S}_{d}$, an integer $1\leq p\leq n$ and tensors $T=u_{1}\otimes\cdots\otimes u_{d}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $S=v_{1}\otimes\cdots\otimes v_{d}\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, $F=f_{1}\otimes\cdots\otimes f_{p}\in(\mathbb{K}^{n_{\pi(1)}})^{\ast}\otimes\cdots\otimes(\mathbb{K}^{n_{\pi(p)}})^{\ast}$, the Kronecker product of $T$ and $S$ is
$T\boxtimes S=(u_{1}\otimes v_{1})\otimes\cdots\otimes(u_{d}\otimes v_{d})\in\mathbb{K}^{n_{1}m_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}m_{d}}.$
We denote by $\langle T,F\rangle$ or $\langle F,T\rangle$ the tensor in $\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{\pi(n)}$ obtained by contracting $T$ with $S$:
@@ 2. Notations
$T\oplus S\coloneqq(u_{1}\oplus v_{1})\otimes\cdots\otimes(u_{d}\oplus v_{d})\in\mathbb{F}_{q}^{n_{1}+m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}+m_{d}}.$
We extend these operations for general tensors by linearity.
2.1. Ranks of tensors
For $S\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, if there exist some partition $\pi\in\mathfrak{S}_{d}$, integer $1\leq p\leq d$, tensors $S_{1}\in\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}$ and $S_{2}\in\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(d)}}$ such that
@@ 2.1. Ranks of tensors
$\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\simeq\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}\otimes\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(d)}}.$
Given $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we denote
@@
If $S\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$ can be written as $S=(g_{1}\otimes\cdots\otimes g_{d})T$ for some $g_{i}\in\operatorname{Hom}(\mathbb{K}^{n_{1}},\mathbb{K}^{m_{1}})$, $1\leq i\leq d$, then we write $S\preceq T$. We define cp-rank and subrank of $T$ respectively by
@@
Here $\operatorname{Id}_{n}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ denotes the identity tensor for each positive integer $n\leq\min\{n_{j}\}_{j=1}^{d}$.
@@
$Z_{k}\left(T\right)\coloneqq\left\{(f_{1},\dots,f_{d-1})\in\bigoplus_{1\leq i\neq k\leq d}\left(\mathbb{F}_{q}^{n_{i}}\right)^{\ast}:\langle T,f_{1}\otimes\cdots\otimes f_{d-1}\rangle=0\right\}.$ (5)
@@
$\operatorname{AR}_{\mathbb{K}}(T)\coloneqq-\log_{|\mathbb{K}|}\sum_{(f_{1},\dots,f_{d})\in\bigoplus_{1\leq i\leq d}\left(\mathbb{F}_{q}^{n_{i}}\right)^{\ast}}\chi(\langle T,f_{1}\otimes\cdots\otimes f_{d}\rangle),$ (6)

### v4

@@ Abstract
[…]
1 Introduction
Given a tensor $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, we denote by $\operatorname{PR}_{\mathbb{K}}(T)$ (resp. $\operatorname{SR}_{\mathbb{K}}(T)$, $\operatorname{AR}_{\mathbb{K}}(T)$, $\operatorname{GR}_{\mathbb{K}}(T)$) the partition rank [42] (resp. slice rank [49], analytic rank [25], geometric rank [35]) of $T$ (cf.  (3)–(6)). These ranks were first introduced to study various problems in additive combinatorics [26, 27, 25, 49, 16, 20, 42], but very soon they also received great attention from other fields such as algebraic geometry [34, 22], coding theory [30, 7] and algebraic complexity theory [8, 35]. It is worth noting that the partition rank is equivalent to the Schmidt rank, as tensors can be viewed as multilinear polynomials. The notion of Schmidt rank was first introduced to explore geometric and arithmetic properties of homogeneous polynomials in [43]. Later, it was extended to inhomogeneous polynomials [33, 1]. The strength of polynomials is an analogue of the partition rank, proposed in [2] to study invariants of polynomial ideals. Both strength and Schmidt rank have played central roles in understanding the behaviour of polynomials [27, 28, 50, 39, 18, 2, 6, 31, 32].
Arguably, the most fundamental problems concerning these ranks are their relations to each other and their stability under field extensions. The stability of a function $f$ of tensors under field extensions means the existence of two constants $c\coloneqq c(d,\mathbb{K})>0$ and $C\coloneqq C(d,\mathbb{K})>0$ such that for any $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ and any (finite) algebraic extension $\mathbb{F}/\mathbb{K}$, we have
@@ 1 Introduction
where $T^{\mathbb{F}}$ is the tensor in $\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$ determined by $T$. Here and in the rest of the paper, we use $C(d,\mathbb{K})$ to indicate the dependence of the constant $C$ on $d$ and $\mathbb{K}$. Furthermore, when two tensor functions $f$ and $g$ satisfy $cf\leq g\leq Cf$ for some positive constants $c$ and $C$, we denote $f\asymp g$.
We remark that in the context of algebraic geometry, the stability of $f$ may be rephrased as the invariance of the order of the magnitude of $f$ under base changes.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
We remark that the original formulations of Conjectures 1.11.3 actually require the constant $C$ to be only dependent on $d$. However, for instance, it was proved in [13] that the original version of Conjecture 1.1 is true over finite fields of sufficiently large cardinality depending on the analytic rank. Thus, to completely resolve the original version of Conjecture 1.1, one needs to remove the dependence of the result in [13] on the analytic rank and prove Conjecture 1.1 for fields of small cardinalities. We also want to mention that according to [48, Remark 4], algebraic problems over fields of small cardinalities are usually considered more difficult than those over sufficiently large fields.
@@
According to [1, Theorem 1.13], Conjecture 1.3 implies Conjecture 1.1, assuming that Proposition $\text{III}_{\mathbb{C}}$ in [43] is valid for arbitrary algebraically closed fields of finite characteristic. Only very recently, it was realized that this assumption is not necessary [13, 31]. In Theorem 5.7, we will prove that Conjectures 1.11.3 are all equivalent. In fact, Theorem 5.7 proves that Conjectures 1.11.3 are all equivalent to
@@
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
It is noticeable that Conjecture 1.4 is a weaker version of the direct sum conjecture for partition rank [24], which is an analogue of the direct sum conjecture for cp-rank [47, 29, 36], disproved by [44]. In particular, Conjecture 1.4 is true for slice rank [24]. It is worth noting that partition rank coincides with slice rank for $d=3$. Consequently, Theorem 5.7 directly implies the validity of Conjectures 1.11.3 for $d=3$, results that had been established in [12, 1, 17].
One of the main ingredients in the proof of Theorem 5.7 is the stability of the analytic rank, proved in Theorem 4.2. In earlier works such as [40, Proposition 8.2] and [13, Proof of Corollary 1], it was proved that for any tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$ and integers $l\geq 1,d\geq 2$,
$\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)\leq l^{d-2}\operatorname{AR}_{\mathbb{F}_{q}}(T).$
Theorem 4.2 not only improves the coefficient $l^{d-1}$ to a constant depending only on $d$ and $q$, but it also shows that $\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)$ is bounded below by $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, up to a constant independent to $l$. In Theorem 4.3, we show that $\lim_{l\to\infty}\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)=\operatorname{GR}_{\mathbb{F}_{q}}(T)$ , which is an analogue of [35, Theorem 8.1]. As an interesting application of Theorems 4.2 and 4.3, we prove in Proposition 4.4 that $\operatorname{AR}_{\mathbb{F}_{q}}(T)\asymp\operatorname{GR}_{\mathbb{F}_{q}}(T)$. Previously, it was only known that $\operatorname{GR}_{\mathbb{F}_{q}}(T)$ is bounded above by a constant multiple of $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, see [12, Theorem 1] and [1, Proof of Theorem 1.13 (1)]. Meanwhile, [3, Theorem 1.2] independently establishes the linear relation between geometric rank and analytic rank, using tools from number theory and algebraic geometry.
As a corollary of the proof of Theorem 5.7, we obtain Proposition 5.4 concerning the stability of slice rank. Although a stronger stability result [17] may be obtained by geometric invariant theory, our proof of Proposition 5.4 is much more elementary. The slice rank $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ (cf.  (24)) of a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ was introduced in [1, Definition 1.3] and it was conjectured that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is also stable under field extensions. In particular, if $\mathsf{W}=\operatorname{span}\{T\}$, then we have $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T)$ by [1, Claim 1.4].
@@
[1, Conjecture 1.16] There is a constant $C\coloneqq C(d)>0$ such that for any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq C\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes_{\mathbb{K}}\overline{\mathbb{K}}).$
@@
To conclude this section, we summarize the main results of this paper in Figure 1, where $T$ is a tensor and $\mathsf{W}$ is a linear subspace of tensors, $\mathbb{F}$ is a finite extension of $\mathbb{F}_{q}$. The symbol $\asymp$” means equal up to a constant factor, $\asymp?$” means $\asymp$” is conjectured to hold and all conjectures around $\operatorname{PR}_{\mathbb{F}_{q}}(T)$ are equivalent.
@@
Figure 1: Summary of results
2 Notations
We first recall three basic operations. Given a permutation $\pi\in\mathfrak{S}_{d}$, an integer $1\leq p\leq n$ and tensors $T=u_{1}\otimes\cdots\otimes u_{d}\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, $S=v_{1}\otimes\cdots\otimes v_{d}\in\mathbb{K}^{m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{m_{d}}$, $F=f_{1}\otimes\cdots\otimes f_{p}\in(\mathbb{K}^{n_{\pi(1)}})^{\ast}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}(\mathbb{K}^{n_{\pi(p)}})^{\ast}$, the Kronecker product of $T$ and $S$ is
$T\boxtimes_{\mathbb{K}}S=(u_{1}\otimes v_{1})\otimes\cdots\otimes(u_{d}\otimes v_{d})\in\mathbb{K}^{n_{1}m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}m_{d}}.$
We denote by $\langle T,F\rangle$ or $\langle F,T\rangle$ the tensor in $\mathbb{K}^{n_{\pi(p+1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{\pi(n)}$ obtained by contracting $T$ with $F$:
@@ 2 Notations
$T\oplus S\coloneqq(u_{1},0)\otimes\cdots\otimes(u_{d},0)+(0,v_{1})\otimes\cdots\otimes(0,v_{d})\in\mathbb{K}^{n_{1}+m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}+m_{d}},$
where $(u,0)$ (resp. $(0,v)$) denotes the image of $u\in\mathbb{K}^{n}$ (resp. $v\in\mathbb{K}^{m}$) in $\mathbb{K}^{n+m}$ via the natural embedding. We extend these operations for general tensors by linearity.
2.1 Ranks of tensors
For $S\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, if there exists some partition $\pi\in\mathfrak{S}_{d}$, integer $1\leq p\leq d-1$, tensors $S_{1}\in\mathbb{K}^{n_{\pi(1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(p)}}$ and $S_{2}\in\mathbb{K}^{n_{\pi(p+1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(d)}}$ such that
@@ 2.1 Ranks of tensors
$\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}\simeq\mathbb{K}^{n_{\pi(1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(p)}}\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(p+1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(d)}}.$
Given $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, we denote
@@
If $S\in\mathbb{K}^{m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{m_{d}}$ can be written as $S=(g_{1}\otimes\cdots\otimes g_{d})T$ for some $g_{i}\in\operatorname{Hom}(\mathbb{K}^{n_{1}},\mathbb{K}^{m_{1}})$, $1\leq i\leq d$, then we write $S\preceq T$. We define cp-rank and subrank of $T$ respectively by
@@
Here $\operatorname{Id}_{n}\in\mathbb{K}^{n}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n}$ denotes the identity tensor for each positive integer $n\leq\min\{n_{j}\}_{j=1}^{d}$.
@@
$Z_{k}\left(T\right)\coloneqq\left\{(f_{1},\dots,f_{d-1})\in\bigoplus_{1\leq i\neq k\leq d}\left(\mathbb{K}^{n_{i}}\right)^{\ast}:\langle T,f_{1}\otimes\cdots\otimes f_{d-1}\rangle=0\right\}.$ (5)
@@
$\operatorname{AR}_{\mathbb{K}}(T)\coloneqq-\log_{|\mathbb{K}|}\frac{\sum_{(f_{1},\dots,f_{d})\in\bigoplus_{1\leq i\leq d}\left(\mathbb{K}^{n_{i}}\right)^{\ast}}\chi(\langle T,f_{1}\otimes\cdots\otimes f_{d}\rangle)}{|\mathbb{K}|^{n_{1}+\cdots+n_{d}}},$ (6)